The force stretching the D string on a certain guitar is 166 N. The string’s linear mass density is 0.003kg/m. What is the speed of waves on the string?
1
Expert's answer
2020-03-11T14:19:32-0400

The speed of sound waves on the string is given by


"v=\sqrt{\frac{T}{\mu}}=\sqrt{\frac{166\:\rm N}{0.003\:\rm kg/m}}=235\:\rm m/s."